1. Given the following equation:


_____ K2PtCl4 + _____ NH3 g _____ Pt(NH3)2Cl2 + _____ KCl


a) Balance the equation.


b) Determine the theoretical yield of KCl if you start with 34.5 grams of NH3.


c) Starting with 34.5 g of NH3, and you isolate 76.4 g of Pt(NH3)2Cl2, what is the percent yield?


1
Expert's answer
2021-03-11T08:46:26-0500

a) K2PtCl4 + 2NH3 → Pt(NH3)2Cl2 + 2KCl

b) M(NH3) = 17.03 g/mol

n(NH3) "= \frac{m}{M} = \frac{34.5}{17.03} = 2.025 \;mol"

According to the reaction equation:

n(KCl) = n(NH3) = 2.025 mol

M(KCl) = 74.55 g/mol

m(KCl) "= n \times M = 2.025 \times 74.55 = 150.96 \;g"

The theoretical yield of KCl is 150.96 g.

c) n(NH3) "= \frac{m}{M} = \frac{34.5}{17.03} = 2.025 \;mol"

According to the reaction equation:

n(Pt(NH3)2Cl2) = "\frac{1}{2}" n(NH3) = 1.012 mol

M(Pt(NH3)2Cl2) = 301.1 g/mol

m(Pt(NH3)2Cl2) "= n \times M = 1.012 \times 301.1 = 304.71 \;g"

Proportion:

304.71 g – 100 %

76.4 g – x

"x = \frac{76.4 \times 100}{304.71} = 25 \%"

The percent yield is 25 %.